OpenStudy (anonymous):

Solve: sqrtx+1 +4=2x

OpenStudy (anonymous):

\[\sqrt{x} + + 4 = 2x\]

OpenStudy (anonymous):

is it \[\sqrt{x + 1} + 4 = 2x\] ???

OpenStudy (anonymous):

Harkirat, that is correct

OpenStudy (anonymous):

\[\sqrt{x + 1} = 2x -4\]

OpenStudy (anonymous):

Got that

OpenStudy (anonymous):

square both sides x + 1 = (2x - 4)² x + 1 = 4x² - 16x + 16 4x² - 16x - x + 16 - 1 = 0 4x² - 17x + 15 = 0 Can u solve the quadratic equation now ?????

OpenStudy (anonymous):

4x² - 17x + 15 = 0 4x² -12x - 5x + 15 = 0 4x(x -3) -5(x - 3) = 0 (4x - 5)(x - 3) = 0 this can be true only if either 4x-5 = 0 or x-3 = 0 if 4x - 5 = 0 ==> x = 5/4 if x - 3 = 0 ==> x = 3